# Twists, Humps, and Pebbles: Multilingual Speech Recognition Models Exhibit Gender Performance Gaps

### v2

@@ Abstract
Current automatic speech recognition (ASR) models are designed to be used across many languages and tasks without substantial changes. However, this broad language coverage hides performance gaps within languages, for example, across genders. Our study systematically evaluates the performance of two widely used multilingual ASR models on three datasets, encompassing 19 languages from eight language families and two speaking conditions. Our findings reveal clear gender disparities, with the advantaged group varying across languages and models. Surprisingly, those gaps are not explained by acoustic or lexical properties. However, probing internal model states reveals a correlation with gendered performance gap. I.e., the easier it is to distinguish speaker gender in a language using probes, the more the gap reduces, favoring female speakers. Our results show that gender disparities persist even in state-of-the-art models. Our findings have implications for the improvement of multilingual ASR systems, underscoring the importance of accessibility to training data and nuanced evaluation to predict and mitigate gender gaps. We release all code and artifacts at https://github.com/g8a9/multilingual-asr-gender-gap.
@@
Giuseppe Attanasio, Beatrice Savoldi, Dennis Fucci, Dirk Hovy  Bocconi University, Milan, Italy  University of Trento, Trento, Italy  Fondazione Bruno Kessler, Trento, Italy giuseppe.attanasio3@unibocconi.it
@@ 1 Introduction
Figure 1: Study overview. We transcribe three speech datasets across 19 languages and compute gender performance gaps. Next, investigate the models for possible causes of these gaps.
A new class of multi-task, multilingual neural networks (Radford et al., 2022; Communication et al., 2023; Chu et al., 2023) has recently pushed the boundaries of several speech-related tasks, including automatic speech recognition (ASR). As these models offer support to an increasing number of languages at no cost, they have found widespread adoption and have been integrated into applications for the general public, such as real-time voice transcription. [^1] E.g., http://tcrn.ch/4el8Brr However, while the user base expands, one question has yet to be answered: will usability be the same for everyone? In this paper, we seek an answer to this question by studying how systems understand the voices of different genders and across multiple languages. These two axes of analysis are motivated by complementary interests.
@@ 2 Background
Contemporary speech recognition systems process audio and text separately and use common cross-modal interactions—e.g., OpenAI’s Whisper (Radford et al., 2022) is loosely inspired to a standard Transformer Vaswani et al. (2017). Crucially, multitasking and multilinguality come at virtually no cost for the user. With Whisper, for example, special “task” and “language” tokens can be prepended to the decoder input to change its functions. This strategy allows multi-tasking and multilinguality without architectural changes or fine-tuning (see Figure 1, top-right, in Radford et al. (2022)). Similarly, Meta’s SeamlessM4T (Communication et al., 2023) uses a speech encoder for audio and an encoder-decoder transformer for text (Costa-jussà et al., 2022, NLLB).
@@ 6 Probing Gender in ASR Models
Motivated by this line of research, we measure gender extractability in ASR models and ask (Q2) whether and to what extent it explains gender-based performance gaps. To our knowledge, ours is the first study of this kind. [^12] Interpretability for speech models is a relatively new research avenue. Mohebbi et al. (2023) use context-mixing techniques to explain homophone disambiguation; Pastor et al. (2023) mask word-units to explain intent detection models.
@@
Figure 3 shows gender extractability on the English CV split for the F-M setup. Trends are similar across languages (full results in Figure 9(a)). First, F1 scores suggest that gender extractability is relatively easy. Whisper produces representations of female and male speech that can be easily separated. Second, extractability is not constant over time. It starts high in the first milliseconds, drops during the actual signal (i.e., the speaker is talking), and finally plateaus around the initial value. This finding recalls the “attention sink” theory, implicating that transformers prioritize initial positions for specific tasks Xiao et al. (2023). Third, MDL probes share similar trends—low codelength indicates an easier task / higher extractability—suggest that simple logistic regression is a robust method and countering the issues found in NLP (Voita and Titov, 2020). [^13] To confirm this hypothesis, we conduct additional experiments by training logistic probes with randomly shifted labels. We report random guess performance on all languages (see Figure 9(b)) in Appendix B.3.
@@
F1S has a weak linear correlation with $\phi(r_{M})$ and $\phi(r_{F})$ (Pearson’s $\rho$ is 0.13 and -0.003, respectively), hinting that F1S cannot explain per-group quality. However, $\rho$ between F1S and $E(r_{F},r_{M})$ is -0.65 ($n=19$, see Figure 8 in Appendix B.3). In other words, linear correlation suggests that the better probes can extract gender from Whisper’s hidden states, the lower the F-M error rate gap is. Note that, as per its definition, $E(r_{F},r_{F})$ can be negative (see points below the zero line in Figure 8). Negative values indicate performance favoring the minority group (here, women). This finding suggests that Whisper does encode recordings from speakers that identify as men and women differently. This aspect can serve as a proxy for measuring and mitigating gender disparities. However, we caution the reader from attributing high extractability to gendered voice and discourage using probes on people’s voice to predict gender (see Ethical considerations, §Ethical Considerations).
@@ 7 Discussion
Empirical studies like ours rely the quality and quantity of evaluation data. To test speech recognition, we need representative splits from diverse speakers. However, current datasets are gender-skewed, with men often overrepresented in terms of records and speakers. While downsampling can help at evaluation time, it does not address the lack of speaker diversity. Skewed distributions are even more evident in low-resource languages when building standard train/validation/test sets is required—e.g., in CV, Yoruba counts only 14 and 20 unique males and females, respectively. Catalan counts only 8 within “Other.”
Therefore, we call for much needed considerations when comparing gender groups in high-resource languages, and new collective efforts to collect more representative data for low-resource languages and gender identities beyond then binary.
@@ B.3 Probing Analysis
Figure 8: Logistic probe F1S vs. $E(r_{F},r_{M})$. CV. Pearson’s correlation (purple dashed line) is -0.65.

### v3

@@ Abstract
Current automatic speech recognition (ASR) models are designed to be used across many languages and tasks without substantial changes. However, this broad language coverage hides performance gaps within languages, for example, across genders. Our study systematically evaluates the performance of two widely used multilingual ASR models on three datasets, encompassing 19 languages from eight language families and two speaking conditions. Our findings reveal clear gender disparities, with the advantaged group varying across languages and models. Surprisingly, those gaps are not explained by acoustic or lexical properties. However, probing internal model states reveals a correlation with gendered performance gap. That is, the easier it is to distinguish speaker gender in a language using probes, the more the gap reduces, favoring female speakers. Our results show that gender disparities persist even in state-of-the-art models. Our findings have implications for the improvement of multilingual ASR systems, underscoring the importance of accessibility to training data and nuanced evaluation to predict and mitigate gender gaps. We release all code and artifacts at https://github.com/g8a9/multilingual-asr-gender-gap.
@@
Giuseppe Attanasio, Beatrice Savoldi, Dennis Fucci, Dirk Hovy  Instituto de Telecomunicações, Lisbon, Portugal  University of Trento, Trento, Italy  Fondazione Bruno Kessler, Trento, Italy  Bocconi University, Milan, Italy giuseppe.attanasio@lx.it.pt
@@ 1 Introduction
Figure 1: Study overview. We transcribe three speech datasets across 19 languages and compute gender performance gaps. Next, we investigate the models for possible causes of these gaps.
A new class of multi-task, multilingual neural networks (Radford et al., 2022; Communication et al., 2023; Chu et al., 2023) has recently pushed the boundaries of several speech-related tasks, including automatic speech recognition (ASR). As these models offer support to an increasing number of languages at no cost, they have found widespread adoption and have been integrated into applications for the general public, such as real-time voice transcription. [^1] See, for instance, http://tcrn.ch/4el8Brr However, while the user base expands, one question has yet to be answered: will usability be the same for everyone? In this paper, we seek an answer to this question by studying how systems understand the voices of different genders and across multiple languages. These two axes of analysis are motivated by complementary interests.
@@ 2 Background
Contemporary speech recognition systems process audio and text separately and use common cross-modal interactions—e.g., OpenAI’s Whisper (Radford et al., 2022) is loosely inspired by the standard Transformer Vaswani et al. (2017). Crucially, multitasking and multilinguality come at virtually no cost for the user. With Whisper, for example, special “task” and “language” tokens can be prepended to the decoder input to change its functions. This strategy allows multi-tasking and multilinguality without architectural changes or fine-tuning (see Figure 1, top-right, in Radford et al. (2022)). Similarly, Meta’s SeamlessM4T (Communication et al., 2023) uses a speech encoder for audio and an encoder-decoder transformer for text (Costa-jussà et al., 2022, NLLB).
@@ 6 Probing Gender in ASR Models
Motivated by this line of research, we measure gender extractability in ASR models and ask (Q2) whether and to what extent it explains gender-based performance gaps. To our knowledge, ours is the first study of this kind. [^12] Interpretability for speech models is a relatively new research avenue. Mohebbi et al. (2023) use context-mixing techniques to explain homophone disambiguation; Pastor et al. (2023) mask word-units to explain intent detection models. Following our work, Krishnan et al. (2024) use amnesic probing to linearly erase gender information in transformer-based ASR models, examining its effects on downstream performance.
@@
Figure 3 shows gender extractability on the English CV split for the F-M setup. Trends are similar across languages (full results in Figure 9(a)). First, F1 scores suggest that gender extractability is relatively easy. Whisper produces representations of female and male speech that can be easily separated. Second, extractability is not constant over time. It starts high in the first milliseconds, drops during the actual signal (i.e., the speaker is talking), and finally plateaus around the initial value. This finding recalls the “attention sink” theory, implicating that transformers prioritize initial positions for specific tasks Xiao et al. (2023). Third, the trends observed in MDL probes—where lower codelength indicates an easier task and higher extractability—suggest that simple logistic regression is a robust method, countering the issues identified in NLP (Voita and Titov, 2020). [^13] To confirm this hypothesis, we conduct additional experiments by training logistic probes with randomly shifted labels. We report random guess performance on all languages (see Figure 9(b)) in Appendix B.3.
@@
Figure 4: Logistic probe F1S vs. $E(r_{F},r_{M})$. CV. The purple dashed line is a linear interpolation.
F1S has a weak linear correlation with $\phi(r_{M})$ and $\phi(r_{F})$ (Pearson’s $\rho$ is 0.13 and -0.003, respectively), hinting that F1S cannot explain per-group quality. However, $\rho$ between F1S and $E(r_{F},r_{M})$ is -0.65 ($n=19$, see Figure 4). In other words, linear correlation suggests that the better probes can extract gender from Whisper’s hidden states, the lower the F-M error rate gap is. Note that, as per its definition, $E(r_{F},r_{F})$ can be negative (see points below the zero line in Figure 4). Negative values indicate performance favoring the minority group (here, women). This finding suggests that Whisper does encode recordings from speakers that identify as men and women differently. This aspect can serve as a proxy for measuring and mitigating gender disparities, e.g. by reducing gender extractability from hidden representations (Krishnan et al., 2024). However, we caution the reader from attributing high extractability to gendered voice and discourage using probes on people’s voice to predict gender (see Ethical considerations, §Ethical Considerations).
@@ 7 Discussion
Empirical studies like ours rely on the quality and quantity of evaluation data. To test speech recognition, we need representative splits from diverse speakers. However, current datasets are gender-skewed, with men often overrepresented in terms of records and speakers. While downsampling can help at evaluation time, it does not address the lack of speaker diversity. Skewed distributions are even more evident in low-resource languages when building standard train/validation/test sets is required—e.g., in CV, Yoruba counts only 14 and 20 unique males and females, respectively. Catalan counts only 8 within “Other.”
Therefore, we call for much needed considerations when comparing gender groups in high-resource languages, and new collective efforts to collect more representative data for low-resource languages and gender identities beyond the binary.
@@ Methods.
Acknowledgments
We thank the reviewers, the members of the SARDINE and MilaNLP research groups, and Marco Gaido for the insightful comments. Giuseppe Attanasio was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI) and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020. He conducted part of the work as a member of the MilaNLP group at Bocconi University, Milan. Dirk Hovy was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 949944, INTEGRATOR) and a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA). He is director of the Data and Marketing Insights Unit of the Bocconi Institute for Data Science and Analysis (BIDSA). Beatrice Savoldi was supported by the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU.